OpenStudy (gabylovesyou):

find f(g(x)) f(x) = 4x - 7 g(x) = 2x - 5 is it f(g(x)) = 8x - 12 ?

OpenStudy (gabylovesyou):

@SolomonZelman

OpenStudy (anonymous):

g(-5x) = -15x-5

OpenStudy (anonymous):

this is how you get that answer

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \color{red}{f(x)} = \color{red}{4x - 7 } }\) \(\large\color{black}{ \displaystyle \color{blue}{g(x)} = \color{blue}{2x - 5 } }\) \(\large\color{black}{ \displaystyle \color{red}{f(\color{blue}{g(x)})} = \color{red}{4(\color{blue}{2x - 5 })- 7 } }\)

OpenStudy (anonymous):

= f(3x - 5) = -5(3x - 5) = -15x + 25 = 25 - 15x g[f(x)] = g(-5x) = 3(-5x) - 5 = -15x - 5

OpenStudy (solomonzelman):

see how I am plugging the g(x) into the f(x) ?

OpenStudy (solomonzelman):

then you just need to simplify.

OpenStudy (gabylovesyou):

yes what comes after

OpenStudy (gabylovesyou):

8x - 5 - 7 ?

OpenStudy (gabylovesyou):

then 8x - 12

OpenStudy (gabylovesyou):

ok

OpenStudy (solomonzelman):

we will expand the parenthesis. \(\large\color{black}{ \displaystyle \color{red}{f(\color{blue}{g(x)})} = \color{red}{4(\color{blue}{2x - 5 })- 7 } }\) \(\large\color{black}{ \displaystyle \color{red}{f(\color{blue}{g(x)})} = \color{red}{4} \times \color{blue}{2x } +\color{red}{4} \times \color{blue}{(-5) ~~-7 } }\)

OpenStudy (solomonzelman):

Again, I am multiplying everything in the parenthesis times 4. By doing so I am expanding the parenthesis.

OpenStudy (solomonzelman):

can you finish simplifying the left side?

OpenStudy (solomonzelman):

the right** side

OpenStudy (gabylovesyou):

8x + 4 * 35

OpenStudy (gabylovesyou):

8x + 140 ?

OpenStudy (solomonzelman):

wow

OpenStudy (solomonzelman):

you are first multiplying the 4 timers (-5) and then subtracting -7.

OpenStudy (solomonzelman):

*times*

OpenStudy (gabylovesyou):

8x + -20 - 7

OpenStudy (gabylovesyou):

?

OpenStudy (solomonzelman):

yes, you multiplied correctly, now subtract

OpenStudy (solomonzelman):

-20-7=?

OpenStudy (gabylovesyou):

-27

OpenStudy (gabylovesyou):

so the answer is 8x + -27 ?

OpenStudy (solomonzelman):

yes, but you can re-write it better

OpenStudy (solomonzelman):

+-27 is same as just -27

OpenStudy (solomonzelman):

instead of having +- there

OpenStudy (gabylovesyou):

so 8x - 27

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

lagging-:( yw

OpenStudy (gabylovesyou):

what about f^-1(x)

OpenStudy (solomonzelman):

inverse function of this new f(g(x)), or inverse function of the f(x) or/and g(x) ?

OpenStudy (solomonzelman):

you want just inverse of f(x) ?

OpenStudy (gabylovesyou):

yes f(x)

OpenStudy (solomonzelman):

I will give you an abstract example.

OpenStudy (gabylovesyou):

its x + 7 / 4 right ?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

(x+7)/4

OpenStudy (gabylovesyou):

ok then thanks

OpenStudy (solomonzelman):

I guess no need for abstract example:) You got it!